# Reverse Diffusion Sequential Monte Carlo Samplers

### v1

@@ Abstract
We propose a novel sequential Monte Carlo (SMC) method for sampling from unnormalized target distributions based on a reverse denoising diffusion process. While recent diffusion-based samplers simulate the reverse diffusion using approximate score functions, they can suffer from accumulating errors due to time discretization and imperfect score estimation. In this work, we introduce a principled SMC framework that formalizes diffusion-based samplers as proposals while systematically correcting for their biases. The core idea is to construct informative intermediate target distributions that progressively steer the sampling trajectory toward the final target distribution. Although ideal intermediate targets are intractable, we develop exact approximations using quantities from the score estimation-based proposal, without requiring additional model training or inference overhead. The resulting sampler, termed Reverse Diffusion Sequential Monte Carlo, enables consistent sampling and unbiased estimation of the target’s normalization constant under mild conditions. We demonstrate the effectiveness of our method on a range of synthetic targets and real-world Bayesian inference problems.
@@ 1 Introduction
Sampling from unnormalized target distributions is a fundamental problem in many applications, ranging from Bayesian inference [1, 2] to simulating molecular systems [3]. Classical methods like Markov chain Monte Carlo (MCMC) simulate a Markov chain with the target as its stationary distribution, but they can suffer from slow mixing and difficulty traversing between modes for complex distributions. Particle-based methods, such as importance sampling, generate exact samples in a large compute limit; yet they struggle with the curse of dimensionality [4]. Alternatively, variational inference (VI) [5] reframes inference as an optimization task, though its success depends on the expressiveness of the variational family and the complexity of the optimization landscape.
@@
Recently, diffusion models have gained traction as a powerful class of generative models for sampling from complex distributions. Diffusion models construct a forward noising process that gradually transforms a complex target distribution into a simple reference distribution [6, 7]. A reverse denoising process is then used to recover target samples from the reference, guided by a time-dependent score function. In the generative modeling setting, this score function is approximated by a neural network trained on samples from the target distribution.
In the sampling context, where such training data is unavailable, recent works have extended the diffusion paradigm by leveraging information from the unnormalized target density. Huang et al. [8], Grenioux et al. [9] estimate the score function using MCMC samples from denoising posteriors of the target, and then simulate from a discretized reverse diffusion process guided by these estimates. This approach demonstrates both theoretical and empirical advantages over conventional Langevin-style MCMC methods, particularly for multi-modal distributions.
A parallel line of research explores diffusion-based neural samplers, which train a neural network to approximate the score function using only the target information. Phillips et al. [10], Akhound-Sadegh et al. [11], OuYang et al. [12] use novel score matching losses, while VI-based methods [13, 14, 15, 16, 17] learn the score by optimizing a divergence between forward and reverse processes.
@@
While promising, diffusion-based sampling methods suffer from two sources of bias: discretization error in simulating the reverse diffusion dynamics and approximation error in the estimated or learned score function. Exceptions are Phillips et al. [10] and Chen et al. [18], which reduce the bias of diffusion-based neural samplers using Sequential Monte Carlo (SMC) – a general tool for sequential inference [19]. However, a recent survey finds that neural samplers still fall short in training efficiency and often rely on special network preconditioning, compared to classical sampling methods [20].
@@
To address these challenges, we develop a training-free diffusion-based sampler, Reverse Diffusion Sequential Monte Carlo (RDSMC), which offers theoretical guarantees. Inspired by previous work, we formalize the diffusion-based sampler as a proposal mechanism within an SMC framework. At a high level, our algorithm generates multiple trajectories, or particles, from the reverse diffusion process using estimated scores. To correct for estimation and discretization biases, we design intermediate target distributions to guide resampling of particles at each step, progressively steering them toward the final target distribution of interest.
Crucially, our intermediate target distributions are constructed from byproducts of score estimation, without additional training or computational overhead. These targets align with the marginal distributions induced by the final target and the forward noising process, resulting in an exact approximation [21, 22]. In contrast, Phillips et al. [10] learn neural network-based alternatives which only attain the true marginal at the final step, and Chen et al. [18] prescribe an annealing path to guide the particles.
@@
RDSMC belongs to a class of nested SMC methods [23], inheriting the standard SMC guarantees. In particular, it produces asymptotically exact samples from the target in the limit of many particles, and provides an unbiased estimate of the normalization constant for any fixed size of particles.
@@
  • We propose a new SMC algorithm, Reverse Diffusion Sequential Monte Carlo (RDSMC), based on the diffusion paradigm for sampling from an unnormalized distribution.
  • RDSMC is training-free, provides an unbiased estimate of the normalization constant, and is asymptotically exact.
  • We show through experiments that RDSMC outperforms or matches existing diffusion-based samplers and classical geometric annealing-based Annealed Importance Sampling [AIS, 24] and SMC samplers [25] on synthetic targets and Bayesian logistic regression benchmarks.
@@ 2 Background
Diffusion models [6, 7] evolve a complex target distribution $\pi(x)$ into a simple reference distribution $p_{1}(x)$, via a a forward stochastic differential equation (SDE)
$dx(\tau)=f(\tau)x(\tau)d\tau+g(\tau)dB_{\tau},\qquad\tau\in[0,1],$ (1)
where $f(\tau)$ and $g(\tau)$ are the drift and diffusion coefficients, and $B_{\tau}$ is standard Brownian motion.
To generate samples from $\pi(x)$, we simulate a reverse SDE initialized from $x(1)\sim p_{1}(x)$,
$dx(\tau)=\left[f(\tau)x(\tau)-g(\tau)^{2}\nabla_{x}\log p_{\tau}(x)|_{x=x(\tau)}\right]d\tau+g(\tau)d\bar{B}_{\tau},\qquad\tau\in[1,0],$ (2)
where $\nabla_{x}\log p_{\tau}(x)$ is the score function of the marginal density $p_{\tau}(x)$ at time $\tau$, and $\bar{B}_{\tau}$ is the reverse-time Brownian motion.
While the score function is typically intractable, it can be approximated by training a score network on samples from the target $\pi(x)$. Once trained, the score network is substituted into the reverse SDE, which is then discretized to generate samples from $\pi(x)$.
Diffusion-based samplers.
In the sampling setting, we are given access to the unnormalized density of $\pi(x)$ rather than its samples. In this case, score estimation can be translated into a posterior inference problem. For example, the denoising score identity [DSI, 26], or Tweedie’s formula [27], provides an expression for the score function,
$\nabla_{x}\log p_{\tau}(x)$ $=\int\frac{\alpha(\tau)x_{0}-x}{\sigma(\tau)^{2}}p_{\tau}(x_{0}\mid x)dx_{0},$ (3)
where $p_{\tau}(x_{0}\mid x)\propto\pi(x_{0})\mathcal{N}(x\mid\alpha(\tau)x_{0},\sigma(\tau)^{2}\mathbb{I})$ is the posterior with the target as prior and the forward kernel as the likelihood parameterized by $\alpha(\tau)$ and $\sigma(\tau)$ (see §˜A.1 for expressions).
Huang et al. [8], Grenioux et al. [9] estimate the score using DSI via MCMC samples from $p_{\tau}(x_{0}\mid x)$, which is then used to simulate the reverse diffusion process in Eq.˜2, starting from a Langevin-style initialization. See §˜4 for further discussion of related approaches.
@@
Sequential Monte Carlo (SMC) is a particle-based method for approximately sampling from a sequence of distributions defined on variables $x_{0:T}$, terminating at a final distribution of interest. We consider a reverse-time formulation, where SMC evolves a weighted collection of $N$ particles $\{x_{t}^{(i)},w_{t}^{(i)}\}_{i=1}^{N}$ from $t=T$ to 0, gradually approximating the final target.
An SMC sampler requires two key design choices [22], a sequence of intermediate proposals, $q(x_{T})$ and $\{q(x_{t}\mid x_{t+1})\}_{t=0}^{T-1}$, and a sequence of (unnormalized) intermediate target distributions, $\{\gamma_{t}(x_{t:T})\}_{t=0}^{T}$, such that the final target $\gamma_{0}(x_{0:T})$ recovers the distribution of interest.
The procedure initializes with $N$ particles $x_{T}^{(i)}\sim q_{T}(x_{T})$ and weights $w_{T}^{(i)}\leftarrow\gamma_{T}(x_{T}^{(i)})/q(x_{T}^{(i)})$ for $i=1,\cdots,N$. Then for each step $t=T{-}1,\dots,0$ and particle $i=1,\dots,N$, SMC proceeds as follows: (1) resample ancestor $x_{t+1}^{(i)}\sim\text{Multinomial}\big{(}x_{t+1}^{(1:N)},w_{t+1}^{(1:N)}\big{)}$; (2) propagate particle $x_{t}^{(i)}\sim q\big{(}x_{t}\mid x_{t+1}^{(i)}\big{)}$; (3) compute weight $w_{t}^{(i)}\leftarrow\gamma_{t}(x_{t:T}^{(i)})\big{/}\left[\gamma_{t}(x_{t+1:T}^{(i)})q(x_{t}^{(i)}\mid x_{t+1}^{(i)})\right]$.
The final set of weighted particles forms a discrete approximation to the final target $\gamma_{0}$, which is asymptotically exact given infinite particles under regularity conditions [19]. However, the efficacy of SMC greatly depends on the choice of intermediate proposals and targets – the closer these intermediate distributions are to the marginals of the final target, the more effective the SMC sampler.
@@ 3 Method
Our goal is to sample from a target distribution $\pi(x)=\frac{1}{Z}\tilde{\pi}(x)$, where $\tilde{\pi}(x)$ is the unnormalized target and the normalization constant $Z=\int\tilde{\pi}(x)dx$ is typically intractable. We develop Reverse Diffusion Sequential Monte Carlo (RDSMC), a diffusion-based SMC sampler targeting $\pi(x)$.
To establish the connection between diffusion models and SMC, we consider an extended target distribution over the trajectory $x_{0:T}=(x_{0},\dots,x_{T})$ by discretizing the forward process in Eq.˜1,
$\pi(x_{0:T})=\pi(x_{0})\prod_{t=1}^{T}p(x_{t}|x_{t-1}),$ (4)
where $p(x_{t}|x_{t-1})$ represents the forward transition kernel. For convenience, we assume a uniform time discretization $0=\tau_{0}<\cdots<\tau_{T}=1$ with step size $\delta=\frac{1}{T}$, and set each $x_{t}:=x(\tau_{t})$. We also write $\alpha_{t}:=\alpha(\tau_{t}),\sigma_{t}:=\sigma(\tau_{t}),f_{t}:=f(\tau_{t}),g_{t}:=g(\tau_{t})$ for time-dependent coefficients, and denote the marginal at step $t$ by $p(x_{t}):=p_{\tau_{t}}(x(\tau_{t}))$ and the posterior by $p(x_{0}\mid x_{t}):=p_{\tau}(x(0)\mid x(\tau_{t}))$.
The extended target $\pi(x_{0:T})$ admits a sequential factorization and preserves the desired marginal $\pi(x_{0})$ at the final step $t=0$, enabling the application of SMC. In §˜3.1, we first introduce a reverse diffusion proposal that approximately samples from the extended target. Building on prior work, the proposal simulates a reverse diffusion process using Monte Carlo (MC) estimates of the score, which introduce auxiliary variables $\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}$} to represent randomness in estimates. In §˜3.2, we construct informative intermediate targets over all variables $\{x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\}$ to correct for proposal bias. These targets account for auxiliary randomness, and approximate the ideal ones using byproducts of the score estimation. Finally in §˜3.3, we present the full RDSMC algorithm and its theoretical guarantees.
@@ 3.1 Reverse diffusion proposal
We devise a sequence of intermediate proposal distributions based on the discretized reverse diffusion process using MC estimates of the score. A crucial component is the explicit treatment of the auxiliary randomness introduced by score estimation, which we model jointly with the diffusion dynamics. The resulting proposal is an extended distribution over both diffusion states and auxiliary variables.
Monte Carlo score estimation.
We first define a generic MC estimator $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ of the score function based on the denoising score identity in Eq.˜3, where $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} denotes the randomness in estimation.
As an illustrative example, consider an importance sampling (IS) estimator (see Algorithm˜2). Given $x_{t}$, we use a proposal $\mathcal{N}(\cdot\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2}\mathbb{I})$ obtained by reversing the forward transition kernel. The score can then be estimated by drawing $M$ samples $\color[rgb]{0.3125,0.46875,0.62890625}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}=\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{(1)},\cdots,{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{(M)}\}$} from this proposal:
$\nabla_{x_{t}}\log p(x_{t})$ $\approx s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}):=\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}},$ (5)
where the importance weights are given by
$w^{(m)}=\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})\,\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}},\sigma_{t}^{2}I)}{\mathcal{N}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2}\mathbb{I})},\qquad m=1,\cdots,M.$ (6)
More informative IS proposals or advanced MC methods such as AIS or SMC can be used to improve score estimates. Other score identities may also be leveraged. See §˜B.1 for further details.
For generality, we define $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} as the collection of all auxiliary random variables generated during the MC score estimation at step $t$, with sampling distribution $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})$. The density $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})$ need not be evaluated. We denote the resulting score estimator as $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
Discretizing reverse diffusion.
With the score estimator in place, we approximate the discretized transition kernel of the reverse SDE in Eq.˜2 at step $t$ by
$q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ $:=\mathcal{N}(x_{t}\mid x_{t+1}-\left[f_{t+1}x_{t+1}-g_{t+1}^{2}s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\right]\delta,g_{t+1}^{2}\delta).$ (7)
@@
The full sampling process begins with drawing $x_{T}$ from the reference distribution $q(x_{T}):=p(x_{T})$, followed by iteratively refining $x_{t}$ through the reverse kernel $q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ in Eq.˜7. At each iteration, the auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\sim q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})$ is introduced to estimate the score $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$. This process defines a proposal over the extended space of $\{x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\}_{t=0}^{T}$,
$q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}})$ $=q(x_{T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})\prod_{t=0}^{T-1}q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}).$ (8)

### v2

@@ Abstract
We propose a novel sequential Monte Carlo (SMC) method for sampling from unnormalized target distributions based on a reverse denoising diffusion process. While recent diffusion-based samplers simulate the reverse diffusion using approximate score functions, they can suffer from accumulating errors due to time discretization and imperfect score estimation. In this work, we introduce a principled SMC framework that formalizes diffusion-based samplers as proposals while systematically correcting for their biases. The core idea is to construct informative intermediate target distributions that progressively steer the sampling trajectory toward the final target distribution. Although ideal intermediate targets are intractable, we develop exact approximations using quantities from the score estimation-based proposal, without requiring additional model training or inference overhead. The resulting sampler, termed Reverse Diffusion Sequential Monte Carlo, enables consistent sampling and unbiased estimation of the target’s normalization constant under mild conditions. We demonstrate the effectiveness of our method on a range of synthetic targets and real-world Bayesian inference problems.  [^1] Our code is available at https://github.com/LuhuanWu/RDSMC.
@@ 1 Introduction
Sampling from unnormalized target distributions is a fundamental problem in many applications, ranging from Bayesian inference [andrieu2003introduction, robert2004metropolis] to simulating molecular systems [frenkel2000molecular]. Classical methods like Markov chain Monte Carlo (MCMC) simulate a Markov chain with the target as its stationary distribution, but they can suffer from slow mixing and difficulty traversing between modes for complex distributions. Particle methods such as importance sampling generate exact samples in a large compute limit; yet they struggle with the curse of dimensionality [chatterjee2018sample]. Alternatively, variational inference (VI) [blei2017variational] casts inference as an optimization task, though its success depends on the expressiveness of the variational family and the complexity of the optimization landscape.
@@
Recently, diffusion models have emerged as a powerful approach for sampling from complex distributions [ho2020denoising, song2020score]. They define a forward noising process that gradually transforms a complex target distribution into a simple base distribution. A reverse denoising process then reconstructs target samples by simulating the dynamics backward in time, starting from the base distribution and guided by a time-dependent score function. In the generative modeling setting, this score function is approximated by a neural network trained on samples from the target distribution. However, in the sampling context, such training data are unavailable and only an unnormalized target density is accessible.
Recent works on diffusion-based samplers explore alternative ways to approximate the score function directly from the target density, enabling sampling without access to training data. One line of works, known as diffusion Monte Carlo (MC) samplers, estimates the score function using MC methods. huang2023reverse, grenioux2024stochastic consider Langevin-style MC algorithms, but they rely on a good initialization of the reverse diffusion process to ensure theoretical guarantees. he2024zeroth propose an alternative scheme based on rejection sampling which relaxes prior assumptions and improves sampling efficiency in low-dimensional regimes. These approaches demonstrate both theoretical and empirical advantages over conventional MCMC methods, particularly for multi-modal distributions.
In contrast to relying on MC estimation during sampling time, a complementary line of research trains a neural network in advance to approximate the score function. To this end, some works propose new score matching objectives [phillips2024particle, akhound2024iterated, ouyang2024bnem], while variational approaches instead optimize divergences between forward and reverse diffusion processes [zhang2021path, berner2022optimal, richter2023improved, vargas2023denoising, vargas2023transport].
@@
While promising, diffusion-based samplers suffer from two sources of bias: discretization error in simulating the reverse diffusion process and approximation error in the estimated or learned score function. An exception is phillips2024particle which mitigates the bias of a trained diffusion-based sampler using Sequential Monte Carlo (SMC), a general inference tool for sequential models [naesseth2019elements, chopin2020introduction]. However, such training-based methods remain computationally complex compared to classical sampling methods and often rely on special neural network preconditioning [he2025no].
@@
To address these challenges, we develop a new diffusion-based sampler, Reverse Diffusion Sequential Monte Carlo (RDSMC) for sampling from unnormalized target distributions, which is training-free and admits theoretical guarantees. Inspired by prior work, we formalize diffusion MC samplers as proposal mechanisms within an SMC framework. At a high level, RDSMC generates multiple particles from the reverse diffusion dynamics using MC-based score estimates. To correct the bias in proposals, we introduce intermediate target distributions that guide resampling of particles at each step, progressively steering them toward the final target distribution of interest.
Crucially, our intermediate targets are efficiently computed using byproducts of MC-based score estimates, incurring no additional cost. Moreover, they form an exact approximation to the ideal intermediate targets that maximize sampling efficiency, defined by the marginal distributions of an extended final target [naesseth2019elements, andrieu2010particle]. This design helps particles stay closely aligned with the final target throughout the sampling process. In contrast, phillips2024particle learn neural network-based surrogates that require additional training and attain the ideal target only at the final step.
@@
RDSMC belongs to a class of nested SMC methods [naesseth2015nested, naesseth2019high], inheriting the standard SMC guarantees. In particular, it produces asymptotically exact samples from the target in the limit of many particles, and provides an unbiased estimate of the normalization constant for any fixed size of particles.
@@
  • We propose a new SMC algorithm, Reverse Diffusion Sequential Monte Carlo (RDSMC), based on the reverse diffusion process for sampling from unnormalized distributions.
  • RDSMC is training-free and extends existing diffusion MC samplers to achieve asymptotically exact sampling and provide unbiased estimates of the normalization constant, while incurring almost no computational overhead given the same number of final samples.
  • Empirically, RDSMC outperforms or matches existing diffusion MC samplers and classical geometric annealing-based Annealed Importance Sampling (AIS) [neal2001annealed] and SMC samplers [del2006sequential] on synthetic targets and Bayesian logistic regression benchmarks.
@@ 2 Background
Diffusion models [ho2020denoising, song2020score] evolve a complex target distribution $\pi(x)$ into a simple base distribution $\pi_{1}(x)$, e.g. $\pi_{1}(x)=\mathcal{N}(0,1)$, via a forward stochastic differential equation (SDE)
$dx_{t}=f(t)x_{t}dt+g(t)dB_{t},\qquad t:0\to 1,$ (1)
where $f(t)$ and $g(t)$ are the drift and diffusion coefficients, and $B_{t}$ is the standard Brownian motion.
To generate samples from $\pi(x)$, we simulate a reverse SDE initialized from $x_{1}\sim\pi_{1}(x)$,
$dx_{t}$ $=\left[f(t)x_{t}-g(t)^{2}\nabla_{x_{t}}\log\pi_{t}(x_{t})\right]dt+g(t)d\bar{B}_{t},\qquad t:1\to 0,$ (2)
where $\nabla_{x_{t}}\log\pi_{t}(x_{t})$ is the score function of the marginal density $\pi_{t}(x_{t})$ at time $t$ induced by the forward process in Eq.˜1, and $\bar{B}_{t}$ is the reverse-time Brownian motion.
Diffusion MC samplers.
Building on the diffusion model paradigm, recent works use MC methods to simulate the reverse dynamics in Eq.˜2 whose terminal distribution corresponds to the target distribution $\pi\propto\tilde{\pi}$ of interest [huang2023reverse, grenioux2024stochastic, he2024zeroth]. While the score function $\nabla_{x_{t}}\log\pi_{t}(x_{t})$ is generally intractable, the key idea is to construct MC estimates via the denoising score identity [DSI, vincent2011connection, efron2011tweedie],
$\nabla_{x_{t}}\log\pi_{t}(x_{t})$ $=\int\frac{\alpha(t)x_{0}-x_{t}}{\sigma(t)^{2}}\pi(x_{0}\mid x_{t})dx_{0},$ (3)
where $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{0}\mid\alpha(t)x_{0},\sigma(t)^{2}\mathbb{I})$ is the denoising posterior, which treats the unnormalized target $\tilde{\pi}(x_{0})$ as the prior and the forward transition density $\mathcal{N}(x_{0}\mid\alpha(t)x_{0},\sigma(t)^{2}\mathbb{I})$ from Eq.˜1 as the likelihood. The coefficients $\alpha(t)$ and $\sigma(t)$ are determined by the drift and diffusion terms $f(t)$ and $g(t)$ (see §˜A.1 for details).
Eq.˜3 suggests that score estimation can be cast as a posterior inference problem. In practice, one can draw approximate samples from the denoising posterior to form an MC estimate of the score, which is then substituted into the reverse dynamics to generate samples from $\pi$. Other score identities beyond the DSI can also be leveraged (see §˜A.2).
@@
SMC [naesseth2019elements, chopin2020introduction] is a particle-based method for sampling from a sequence of distributions defined on variables $x_{0:T}$, terminating at a final target distribution of interest. We consider a reverse-time formulation, where SMC evolves a weighted collection of $N$ particles $\{x_{t}^{(i)},w_{t}^{(i)}\}_{i=1}^{N}$ from $t=T$ to 0, gradually approximating the final target.
An SMC sampler requires two key design choices [naesseth2019elements], a sequence of intermediate proposals, $q_{T}(x_{T})$ and $\{q_{t}(x_{t}\mid x_{t+1})\}_{t=0}^{T-1}$, and a sequence of (unnormalized) intermediate target distributions, $\{\gamma_{t}(x_{t:T})\}_{t=0}^{T}$, such that the final target $\gamma_{0}(x_{0:T})$ recovers the distribution of interest.
SMC initializes $N$ particles $x_{T}^{(i)}\sim q_{T}(x_{T})$ with weights $w_{T}^{(i)}\leftarrow\gamma_{T}(x_{T}^{(i)})/q_{T}(x_{T}^{(i)})$ for $i=1,\cdots,N$. Then for each step $t=T{-}1,\dots,0$ and particle $i=1,\dots,N$, it proceeds as follows:
  1. 1.
    resample ancestor $x_{t+1}^{(i)}\sim\text{Multinomial}\big(x_{t+1}^{(1:N)},w_{t+1}^{(1:N)}\big)$;
  2. 2.
    propagate particle $x_{t}^{(i)}\sim q_{t}\big(x_{t}\mid x_{t+1}^{(i)}\big)$;
  3. 3.
    compute weight $w_{t}^{(i)}\leftarrow\gamma_{t}(x_{t:T}^{(i)})\big/\left[\gamma_{+1}(x_{t+1:T}^{(i)})q_{t}(x_{t}^{(i)}\mid x_{t+1}^{(i)})\right]$.
The final set of weighted particles forms a discrete approximation to the final target $\gamma_{0}$, which is asymptotically exact given infinite particles under regularity conditions [chopin2020introduction]. However, the efficacy of SMC greatly depends on the choice of intermediate targets and proposals – the closer these intermediate distributions are to the marginals of the final target, the more effective the SMC sampler.
@@ 3 Method
Our goal is to sample from a target distribution $\pi(x)=\frac{1}{Z}\tilde{\pi}(x)$, where $\tilde{\pi}(x)$ is the unnormalized target and $Z=\int\tilde{\pi}(x)\mathrm{d}x$ is a generally intractable normalization constant. We develop Reverse Diffusion Sequential Monte Carlo (RDSMC), a diffusion-based SMC sampler targeting $\pi(x)$.
To enable sequential inference with SMC, we define an extended target distribution over a discretized trajectory $x_{0:T}=(x_{0},\dots,x_{T})$ of the forward process in Eq.˜1,
$\pi(x_{0:T})=\pi(x_{0})\prod_{t=1}^{T}\pi(x_{t}|x_{t-1}),$ (4)
where $0=\tau_{0}<\cdots<\tau_{T}=1$ are $T+1$ discretization times and $\pi(x_{t}|x_{t-1})$ is the forward transition density from time $\tau_{t-1}$ to $\tau_{t}$ induced by Eq.˜1 (see §˜A.1 for analytical expressions). Without loss of generality, we assume a uniform discretization step size $\delta=1/T=\tau_{t}-\tau_{t-1},\forall t=1,\cdots,T$.
This construction yields a sequence of intermediate targets $\{\pi(x_{t:T})\}_{t=T}^{0}$, whose final marginal $\pi(x_{0})$ recovers the desired target. These intermediate targets provide a natural setting for SMC, and in fact, are optimal intermediate gargets. However, they are intractable to sample from and to evaluate.
To address this challenge, we develop RDSMC leveraging the dual structure of diffusion processes, generating samples in the reverse direction while grounding their targets in the forward direction. In §˜3.1, we introduce a sequence of proposals based on the reverse diffusion process using MC score estimates. To correct the resulting proposal bias, in §˜3.2, we develop practical intermediate targets that form exact approximations to the optimal targets $\{\pi(x_{t:T})\}_{t=0}^{T}$ using byproducts of MC score estimates. Finally in §˜3.3, we present the full RDSMC algorithm and its theoretical guarantees.
Notation.
We slightly abuse notation by letting $x_{t}$ denote both the continuous-time variable $x_{t}$ for $t\in[0,1]$, as used in Eqs.˜1 and 2, and the discrete-time variable for $t\in\{0,\dots,T\}$, as used throughout this section. We denote $f_{t}\coloneqq f(\tau_{t}),g_{t}\coloneqq g(\tau_{t}),\alpha_{t}\coloneqq\alpha(\tau_{t}),\sigma_{t}\coloneqq\sigma(\tau_{t})$ for time-dependent diffusion coefficients from Eqs.˜1 and 3. For generality, we define $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} as the collection of all auxiliary random variables generated in the MC score estimation at step $t$.
@@ 3.1 Reverse diffusion proposal
We design an extended proposal distribution based on the reverse diffusion process with MC score estimates, jointly modeling diffusion dynamics $x_{t}$ and auxiliary randomness $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} from score estimation.
MC score estimation.
Given a sample $x_{t}$ at step $t$, we define a generic MC score estimator $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\approx\nabla_{x_{t}}\log\pi(x_{t})$ based on the DSI in Eq.˜3, where $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} is the estimation randomness associated with some sampling distribution $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})$.
As an illustrative example, consider an importance sampling (IS) estimator (Algorithm˜2) targeting the posterior $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$. We draw $M$ importance samples $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}$} from the proposal $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\mid x_{t})\coloneqq\mathcal{N}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2}\mathbb{I})$ for $m=1,\cdots,M$, and estimate the score by
$\nabla_{x_{t}}\log\pi(x_{t})$ $\approx s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\coloneqq\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}},$ (5)
which corresponds to an MC approximation of the RHS of Eq.˜3. The importance weights $w^{(m)}$ are
$w^{(m)}=\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})\,\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}},\sigma_{t}^{2}\mathbb{I})}{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\mid x_{t})},\qquad m=1,\cdots,M.$ (6)
In this case, the auxiliary randomness is to the collection of all importance samples ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}=\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\}_{m=1}^{M}$.
In practice, we use a more sophisticated AIS scheme to improve the accuracy of score estimates (Algorithm˜3). More informative IS proposals or other MC methods such as SMC and rejection sampling may also be used; see §˜B.1 for further details.
Approximate reverse diffusion dynamics.
With the score estimates in place, we approximate the transition kernel of the reverse SDE in Eq.˜2 to define a conditional proposal for $x_{t}$,
$q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ $\coloneqq\mathcal{N}(x_{t}\mid x_{t+1}-\left[f_{t+1}x_{t+1}-g_{t+1}^{2}s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\right]\delta,g_{t+1}^{2}\delta).$ (7)
@@
The full sampling process starts by drawing $x_{T}$ from a tractable base distribution $q(x_{T})$, e.g. $q(x_{T})=\mathcal{N}(0,1)$. We then iteratively evolve $x_{t}$ for $t=T-1,\cdots,0$ using the reverse diffusion dynamics $q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ in Eq.˜7 where auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}\sim q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}\mid x_{t+1})$ is sampled in previous iteration to estimate the score $s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$.
The overall sampling process defines a proposal over the extended space of $\{x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\}_{t=0}^{T}$,
$q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}})$ $\coloneqq q(x_{T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})\prod_{t=0}^{T-1}q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}).$ (8)